# Gauge-dependence of Scalar Induced Gravitational Waves

### v1

@@ 4 Gauge-dependence of first-order metric perturbations
We now review how quantities in the metric change under coordinate transformation, also known as gauge-transformation. There are two approaches in studying gauge-transformation, the active and passive approach. We follow the approach taken by [33, 34, 18], which calculates the change in perturbations under coordinate change at a fixed physical point in spacetime. Later we look at the construction of the gauge-invariant variables.
@@ 4.1 First-order coordinate transformation of cosmological metric perturbations
where fixing $\xi^{0}$ determines the time-slicing and the spatial coordinate by $\xi$. With these transformations the perturbed line element, Section 2, can be written in terms of the transformed coordinate system. The scalar modes transform as
@@
As mentioned in Section 2, first-order vector modes and tensor modes have negligible contribution to the second-order tensor modes and therefore not considered.
@@ 4.2 Gauge-invariant combination of first-order scalar perturbations
The gauge-invariant approach implies that quantities that are gauge dependent, in this case the scalar perturbations, can be constructed into gauge-invariant variables by defining a combination of perturbations variables which coincide with a particular gauge. This is obtained by fixing the coordinate transformations Eqs. 35, 36, 37 and 38. This is different from the gauge-independent quantities such as the first-order tensor perturbation, which is the same in all gauges. For that reason, one can find different gauge-invariant variables that correspond to different gauge choices and using these variables, the governing equations can be reformulated. Various gauges are chosen depending on the case which is being studied [35]. In the following section, we review gauges that have been used in literature in studying SIGWs.
@@ 4.2.1 Synchronous gauge
which allows us to set $\phi^{(1)}=\hat{w}_{i}^{(1)}=0$. This provides a simplification in the dynamical equations, as the cosmic time coordinates of FRW background coincide with the proper time of a observer at a fixed spatial coordinate, i.e. $d\tau=ad\tau$. One disadvantage with this gauge is that it is not possible to formulate a clear gauge-invariant variable to due residual gauge freedom. The gauge-transformation Eqs. 35 and 37 give the following
@@ 5 Source function of SIGWs in different gauges
In this section, we analyze the scalar-scalar source term of SIGWs in the gauges mentioned above using two approaches: (i) deriving the source term in the synchronous and Poisson gauges via their gauge conditions, and (ii) employing gauge-invariant variables to study the source term in the uniform curvature gauge.
@@ 5.1 Scalar-scalar source in synchronous gauge
Using the definition for the shear potential, $\sigma^{(1)}$, and the constraint $\phi^{(1)}=w^{(1)}=0$, Appendix E simplifies to,
@@ 5.2 Scalar-scalar source in Poisson gauge
By using the gauge-condition $w^{(1)}=h^{(1)}=0$ and in the absence of anisotropic stress where $\phi^{(1)}=\psi^{(1)}$, we can simplify Appendix E in the Poisson gauge
@@ 5.3 Scalar-scalar source in uniform curvature gauge
Next we will study the uniform curvature gauge where $\psi^{(1)}=h^{(1)}=0$. We can reduce Appendix E to
@@ 6 Kernel in various gauges
In this section we derive the analytical expressions of the source function and later perform numerical integration of the kernel in the various gauges. As seen in Section 5, the source function in the synchronous gauge, Section 5.1, is found in terms of the transfer functions for the scalar modes $T_{\psi}$ and $T_{\sigma}$. We study the behavior in synchronous gauge, following the method outlined in [27], by deriving the linear solutions for the scalar modes and subsequently simplify the source term. On the other hand, the source function in the Poisson gauge, Section 5.2, and uniform gauge, Section 5.3, are written in terms of the transfer function for the scalar potential $\phi^{(1)}$. We can repeat the same analysis of the first-order scalar potential in the Poisson gauge. Using the transfer function, $T_{\phi}$, we are able to compute both in the Poisson and uniform gauge.
@@ 6.4 Numerical integration of the kernel
We numerically evaluate the kernel, Eq. 29, in the three different gauges using the source functions Section 6.3, Section 6.3 and Section 6.3. By fixing a set of values for $u$ and $v$, the numerical integration is presented in Fig. 1. We show the comparison of the numerically integrated kernel in all three gauges with the oscillation average of the kernel function, Eq.(26) from [32] which was found in the conformal Newtonian gauge. In addition, we present residual between our numerically integrated kernel and the analytical solution [32].
@@
Figure 1: The kernel function $I^{2}(v=1,u=1,x)$ as a function of $x=k\tau$. In the left panel we show the oscillation average of the kernel function (red line) taken from Eq.(26) of [32] compared to numerical integration of the kernel in the synchronous (green line), Poisson (blue line) and uniform curvature (orange line) gauges. In the right panel, we plot the residual between the numerically integrated kernel in three gauges and the analytical integration.
@@
We observe that the numerical results from the different gauge choices closely follow the analytical solution. The kernel values for the Poisson, uniform curvature, and synchronous gauges exhibit minimal variation, indicating that any residual gauge effects are small. The oscillatory structure seen in the analytic oscillation average seems to align with the numerically integrated kernels. From the panel in the right, we notice the residuals appear small across all gauges, confirming that the gauge choices do not significantly affect the final gravitational wave kernel. The synchronous gauge may introduce this numerical discrepancies due to its residual gauge freedom. Even if there is a minimal discrepancy, the variation in behavior in all three gauges decreases as $x\gg 1$, when entering deep in the horizon. This suggests that despite gauge differences, the kernel remains consistent, confirming gauge invariance at the level of physical observables.
@@ 7 Conclusion
Acknowledgments I would like to thank Sabino Matarrese and Andrea Maselli for their insightful comments and discussions. My gratitude also goes to Marisol Traforetti for her assistance with the computation and numerical integration of the kernels.
@@ D.3 Space - space component of energy-momentum tensor
Appendix E Source term of SIGW
@@
$S^{i}_{j}=$ $\left.\partial^{i}\phi^{(1)}\partial_{j}\phi^{(1)}+2\phi^{(1)}\partial^{i}\partial_{j}\phi^{(1)}-2\psi^{(1)}\partial^{i}\partial_{j}\phi^{(1)}-\partial_{j}\phi^{(1)}\partial^{i}\psi^{(1)}-\partial^{i}\phi^{(1)}\partial_{j}\psi^{(1)}\right.$

### v2

@@ 4 Gauge-dependence of first-order metric perturbations
We now review how quantities in the metric change under coordinate transformation, also known as gauge-transformation. This section clarifies how different choices of time and space coordinates affect the scalar perturbations and introduces gauge-invariant combinations. There are two approaches in studying gauge-transformation, the active and passive approach. We follow the approach taken by [39, 40, 22], which calculates the change in perturbations under coordinate change at a fixed physical point in spacetime. Later we look at the construction of the gauge-invariant variables.
@@ 4.1 First-order coordinate transformation of cosmological metric perturbations
where fixing $\xi^{0}$ determines the time-slicing and the spatial coordinate by $\xi$. With these transformations the perturbed line element can be written in terms of the transformed coordinate system. The scalar modes transform as
@@
As discussed earlier in Section 2, vector and tensor perturbations at first order—$w^{(1)}_{i}$, $h^{(1)}_{i}$, and $h^{(1)}_{ij}$—are neglected due to their minimal contribution to second-order tensor modes.
@@ 4.2 Gauge-invariant combination of first-order scalar perturbations
In cosmological perturbation theory, many quantities—such as scalar perturbations—are inherently gauge-dependent, meaning their values can change under coordinate transformations. The gauge-invariant approach addresses this by constructing combinations of perturbation variables that remain unchanged under such transformations. These combinations are called gauge-invariant variables and are designed to coincide with the values of the perturbations in a specific gauge. This procedure involves fixing the coordinate freedom using transformation rules such as those shown in Eqs. 29, 30, 31 and 32.
It is important to distinguish between gauge-invariant variables and physical observables. Gauge-invariant variables are tools within the theoretical framework that help consistently describe perturbations across different gauges, but they are not necessarily directly measurable. In contrast, observables—such as temperature anisotropies in the cosmic microwave background or gravitational wave signals—are quantities that are invariant under coordinate transformations and correspond to physical measurements. As emphasized in [33, 31], not every gauge-invariant quantity is an observable, though all observables must be gauge-invariant.
Some quantities, like the first-order tensor perturbations, are already gauge-independent—they remain the same in all coordinate systems and are directly related to physical observables like gravitational waves. For gauge-dependent quantities, different gauge choices can lead to different gauge-invariant variables, and the governing equations can be reformulated accordingly. The choice of gauge often depends on the problem being studied and the mathematical convenience it offers [41].
In the following section, we review gauge choices that have been commonly used in the study of scalar-induced gravitational waves (SIGWs) and discuss their implications for theoretical consistency and observational relevance.
@@ 4.2.1 Synchronous gauge
which allows us to set $\phi^{(1)}=\hat{w}{{}_{i}}^{(1)}=0$. This provides a simplification in the dynamical equations, as the cosmic time coordinates of FRW background coincide with the proper time of a observer at a fixed spatial coordinate, i.e. $d\eta=ad\tau$. One disadvantage with this gauge is that it is not possible to formulate a clear gauge-invariant variable to due residual gauge freedom. The gauge-transformation Eqs. 29 and 31 give the following
@@ 5 Source function of SIGWs in different gauges
Here we identify the explicit form of the source term for second-order gravitational waves that arises from the quadratic coupling of first-order scalar modes. This quantity encapsulates the physical mechanism by which density perturbations generate gravitational radiation during horizon re-entry, and its structure is central to SIGW phenomenology. As mentioned in Section 3.3 source term is found in Fourier space as
which is a convolution of Appendix F with primordial fluctuations. We analyze the scalar-scalar source term of SIGWs in the gauges mentioned above using two approaches: (i) deriving the source term in the synchronous and Poisson gauges via their gauge conditions, and (ii) employing gauge-invariant variables to study the source term in the uniform curvature gauge. In all derivations, we retain only first-order scalar perturbations $\phi^{(1)}$, $\psi^{(1)}$, etc., as sources for the second-order tensor perturbation $h^{(2)}_{ij}$.
@@ 5.1 Scalar-scalar source in synchronous gauge
Using the definition for the shear potential, $\sigma^{(1)}$, and the constraint $\phi^{(1)}=w^{(1)}=0$, Appendix F simplifies to,
@@ 5.2 Scalar-scalar source in Poisson gauge
By using the gauge-condition $w^{(1)}=h^{(1)}=0$ and in the absence of anisotropic stress where $\phi^{(1)}=\psi^{(1)}$, we can simplify Appendix F in the Poisson gauge
@@ 5.3 Scalar-scalar source in uniform curvature gauge
Next we will study the uniform curvature gauge where $\psi^{(1)}=h^{(1)}=0$. We can reduce Appendix F to
@@ 6 Kernel in various gauges
In this section we derive the analytical expressions of the source function and later perform numerical integration of the kernel in the various gauges. As seen in Section 5, the source function in the synchronous gauge, Section 5.1, is found in terms of the transfer functions for the scalar modes $T_{\psi}$ and $T_{\sigma}$. We study the behavior in synchronous gauge, following the method outlined in [31], by deriving the linear solutions for the scalar modes and subsequently simplify the source term. On the other hand, the source function in the Poisson gauge, Section 5.2, and uniform gauge, Section 5.3, are written in terms of the transfer function for the scalar potential $\phi^{(1)}$. We can repeat the same analysis of the first-order scalar potential in the Poisson gauge. Using the transfer function, $T_{\phi}$, we are able to compute both in the Poisson and uniform gauge. For the numerical analysis, we use the first-order transfer functions for the scalar perturbations, to derive the analytical expression of the source function in all three gauges. We present the source functions in Appendix G in terms of the transfer functions, where the transfer functions $T_{\psi}(k,\tau)$, $T_{\phi}(k,\tau)$, and $T_{\sigma}(k,\tau)$ describe the time evolution of first-order scalar perturbations in the respective gauges.
@@ 6.3 Numerical integration of the kernel
We numerically evaluate the kernel, Eq. 23, in the three different gauges using the source functions Appendix G, Appendix G and Appendix G. By fixing a set of arbitrary values for $u=1$ and $v=1$, the numerical integration is presented in Fig. 1. The parameters $u$ and $v$ determine the time of horizon entry of the source modes, with higher values corresponding to modes that enter much later. In the radiation-dominated era, the transfer function scales approximately as $q\tau$, and hence becomes suppressed at late times. As a result, configurations with large $u$ and $v$ contribute negligibly to the kernel. To capture the most significant contribution to the gravitational wave signal, we focus primarily on small values of $u$ and $v$, where the sources are most active. We compare the numerically integrated kernel in all three gauges with the analytical squared kernel from Eq. (25) and it’s oscillation-averaged from Eq. (26) of [36], derived in the conformal Newtonian gauge. This is shown in the left panel, where the numerical kernels are compared to the analytical squared kernel. In the right panel, we show the residuals between the numerical squared kernels and the analytical squared kernel from Eq. (25) of [36].
@@
Figure 1: The kernel function $I^{2}(v=1,u=1,x)$ as a function of $x=k\tau$. In the left panel we show the squared analytical kernel (purple line) taken from Eq. (25), and it’s oscillation average (red line) taken from Eq.(26) of [36] compared to numerical integration of the kernel in the synchronous (green line), Poisson (blue line) and uniform curvature (orange line) gauges. In the right panel, we plot the residual between the numerically integrated kernel in three gauges and the analytical integration Eq. (25) from [36].
@@
We observe that the numerical results from the different gauge choices closely follow the analytical solution. The kernel values for the Poisson, uniform curvature, and synchronous gauges exhibit minimal variation, indicating that any residual gauge effects are small. The oscillatory structure seen in the analytical oscillation average aligns well with the numerically integrated kernels. From the right panel, we see that the residuals remain small across all gauges, supporting the conclusion that the choice of gauge does not significantly affect the final gravitational wave kernel. In the region $x\sim 50$ to $x\sim 100$, a modulation in the oscillation amplitude of the numerical kernel becomes apparent, particularly in the synchronous and uniform curvature gauges. This modulation occurs on a scale larger than the individual oscillation period. We interpret this effect as a numerical artefact rather than a physical feature. The modulation is likely related to the numerical integration scheme, which may imperfectly capture the rapidly oscillating integrand at late times.
For larger values of $x$, the synchronous gauge (green curve) appears to lie below the uniform curvature gauge (orange curve), with points of tangency occurring near the oscillation minima of the orange curve. However, this behavior is not systematic and should not be interpreted as a physical lower bound. Upon closer inspection and using higher-resolution plots or narrower vertical limits, the apparent tangency is revealed to be an artifact of visual resolution rather than a genuine feature of the underlying functions. Moreover, the residuals depend on the specific values of the dimensionless variables $u$ and $v$. We illustrate this dependence in Fig. 2, which shows how the residual structure varies significantly with different choices of $u$ and $v$. We also see that the residual between the numerical and analytical results tends to decrease for larger values of $u$ and $v$, as the transfer function becomes increasingly suppressed and the numerical kernel approaches the analytical approximation more closely.
Figure 2: Residuals between the numerically integrated kernels and the analytical expression from Eq. (25) of [36], plotted for different values of the parameters $u$ and $v$. Each subplot corresponds to a different pair of $(u,v)$, illustrating how the residual structure varies with the choice of source mode configuration.
Even if there is a minimal discrepancy, the variation in behavior in all three gauges decreases as $x\gg 1$, when entering deep in the horizon.
6.4 Connecting SIGWs to Present-Day Observables
To assess the observational relevance of scalar-induced gravitational waves (SIGWs), we must relate their spectral energy density at the time of generation during the radiation-dominated (RD) epoch to their present-day energy density spectrum, $\Omega_{\rm GW,0}(k)$. This connection is established under two primary assumptions:
  1. 1.
    The emission of SIGWs occurs after the reheating phase and during the RD epoch, which is when curvature perturbations re-enter the horizon and induce tensor modes.
  2. 2.
    After their generation, the energy density of the GWs redshifts as radiation, $\rho_{\rm GW}\propto a^{-4}$, until matter-radiation equality. Post-equality, this radiation dilution continues in the presence of non-negligible changes in the relativistic degrees of freedom, encoded through the functions $g_{*}$ and $g_{*s}$, representing the effective number of relativistic species and entropy degrees of freedom respectively.
Taking these effects into account, the present-day GW spectral density is given by [44]:
$\Omega_{\rm GW,0}(k)h^{2}=\Omega_{r,0}h^{2}\left(\frac{g_{*s}(T_{\rm hc})}{g_{*s0}}\right)^{-4/3}\left(\frac{g_{*}(T_{\rm hc})}{g_{*0}}\right)\Omega_{\rm GW,hc}(k)\,,$ (58)
where $h=H_{0}/100\,\text{km}\,\text{s}^{-1}\,\text{Mpc}^{-1}$ is the reduced Hubble constant, and the subscript “hc” denotes quantities evaluated at the horizon-crossing time during the RD era. For temperatures $T_{\rm hc}\gtrsim 1\,\mathrm{TeV}$, the Standard Model gives $g_{*}=g_{*s}=106.75$. The current radiation density parameter is constrained by Planck measurements to be $\Omega_{r,0}h^{2}\approx 4.2\times 10^{-5}$ [45].
The only remaining input required to perform the numerical integration of the SIGW spectrum is the primordial curvature power spectrum. Since the SIGW spectrum scales quadratically with the primordial scalar power spectrum, we adopt a log-normal form for the dimensionless spectrum [46]:
$\Delta^{2}_{\zeta}(k)=\frac{\mathcal{A}_{\zeta}}{\sqrt{2\pi\sigma^{2}}}\exp\left(-\frac{\ln^{2}({k/k_{*}})}{2\sigma^{2}}\right)\,,$ (59)
where $\mathcal{A}_{\zeta}$ is the total integrated amplitude, $\sigma$ determines the width of the peak, and $k_{*}$ its position. In our computations, we use $\mathcal{A}_{\zeta}=2.1\times 10^{-2}$ and $\sigma=0.1$, which ensures a narrow peak structure significantly above the CMB amplitude ($A_{\zeta}(k_{\rm CMB})\sim 2.1\times 10^{-9}$ at $k_{\rm CMB}=0.05\,\mathrm{Mpc}^{-1}$), while remaining unconstrained at smaller scales relevant for SIGWs. The peak wavenumber $k_{*}$ can be chosen based on the detector’s sensitivity using the relation $f_{*}=k_{*}/(2\pi)$.
Figure 3: Spectral energy density of scalar-induced gravitational waves (SIGWs), $\Omega_{\rm GW}(k)/\Omega_{r}$, as a function of comoving wavenumber normalized to a reference scale $k/k_{\rm ref}$. Results are shown for three gauge choices: Poisson (blue), synchronous (green), and uniform curvature (orange), computed during the radiation-dominated era.
In Fig. 3, we compare the SIGW energy density spectrum computed in three commonly used gauges—Poisson, synchronous, and uniform curvature. The three curves show good agreement, particularly in the sub-horizon regime ($k\gg k_{\rm ref}$), where the scalar source term becomes negligible and gravitational waves propagate freely. As expected, gauge dependence is most pronounced on super-horizon scales, where the scalar source remains active and its gauge-specific formulation directly affects the evolution of tensor modes. Minor discrepancies at intermediate scales can be attributed to residual gauge effects from numerical integration and the gauge-dependent structure of the source term. These variations diminish at late times, reinforcing the conclusion that the SIGW spectrum is effectively gauge-invariant in the observationally relevant regime. This supports earlier arguments that gauge effects are suppressed once tensor modes decouple from their scalar sources and evolve independently [31, 32].
@@ 7 Conclusion
It is important to emphasize that the approximate gauge invariance we observe applies specifically in the radiation-dominated epoch well inside the horizon ($x\gg 1$) when the scalar sources become inactive. As demonstrated in [32], in a dust-dominated universe or during epochs where sources remain active, residual gauge effects can persist and impact observables. Therefore, the gauge-invariant behavior of SIGWs should be interpreted within the appropriate cosmological context.
@@
Acknowledgments I would like to thank S. Matarrese, A. Maselli and G. Perna for their insightful comments and discussions. My gratitude also goes to M. Traforetti for her assistance with the computation and numerical integration of the kernels.
@@ D.3 Space - space component of energy-momentum tensor
Appendix E Projected spatial Einstein- and energy momentum tensor
Here we present the spatial Einstein tensor Section C.3, and energy-momentum tensor, Section D.3, with the projection tensor, $P^{li}_{jm}$ [34], applied,
$\left.\left.+2\mathcal{H}w^{(1)}+w^{(1)^{\prime}}\right)\right].$ (112)
Appendix F Source term of SIGW
@@
$S^{i}_{{}_{j}}=$ $\left.\partial^{i}\phi^{(1)}\partial_{j}\phi^{(1)}+2\phi^{(1)}\partial^{i}\partial_{j}\phi^{(1)}-2\psi^{(1)}\partial^{i}\partial_{j}\phi^{(1)}-\partial_{j}\phi^{(1)}\partial^{i}\psi^{(1)}-\partial^{i}\phi^{(1)}\partial_{j}\psi^{(1)}\right.$